# Diversity-Aware Policy Optimization for Large Language Model Reasoning

### v1

@@ Appendix A Theoretical analysis
(2). Since $T$ is a random variable, we apply the law of total probability:
@@
Hence, we can apply case (1) and finish the proof.
@@ Appendix B More discussion
We believe the performance gap between our R1-zero reproduction and SimpleRL-Zoo’s implementation stems primarily from resource constraints. Our experiments were conducted on a modest 8×8$\times$8 ×A6000 GPUs setup, necessitating several efficiency optimizations: we employed the simpler GSM8K dataset, constrained generation lengths (appropriate for GSM8K’s short responses). In contrast, SimpleRL-Zoo utilized significantly more powerful 2×8 H100-80G GPUs, trained on more complex datasets with longer response lengths. Importantly, our study’s primary objective was not to surpass SimpleRL-Zoo’s results, but rather to demonstrate that our diversity-enhanced method outperforms standard R1-zero. Our method of independence can be applied to enhance the SimpleRL-Zoo and other state-of-the-art methods.
@@ E.1 Entropy during the RL training
E.2 Solution diversity in R1-zero-Div

### v2

@@ Appendix A Theoretical analysis
(2). For the case that $T$ is a random variable, roughly, the idea is to apply the law of total probability:
@@
And apply case (1) to finish the proof.
@@ Appendix B More discussion
We believe the performance gap between our R1-zero reproduction and SimpleRL-Zoo’s implementation stems primarily from resource constraints. Our experiments were conducted on a modest $8\timesA$6000 GPUs setup, necessitating several efficiency optimizations: we employed the simpler GSM8K dataset, constrained generation lengths (appropriate for GSM8K’s short responses). In contrast, SimpleRL-Zoo utilized significantly more powerful 2×8 H100-80G GPUs, trained on more complex datasets with longer response lengths. Importantly, our study’s primary objective was not to surpass SimpleRL-Zoo’s results, but rather to demonstrate that our diversity-enhanced method outperforms standard R1-zero. Our method of independence can be applied to enhance the SimpleRL-Zoo and other state-of-the-art methods.
@@ E.1 Entropy during the RL training
E.2 Pass@1 Accuracy (on test set) against the training steps
Figure 4: Pass@1 Accuracy (on test set) against the training steps.
E.3 Evaluation on GPQA
We present the avg@8 results on the GPQA Diamond dataset in Table 7, where we directly evaluated the models listed in Table 1. These results demonstrate the generalizability of our method.
Table 7: Avg@8 on GPQA Diamond dataset
Method GPQA Diamond
Qwen2.5-Math-7B 28.5 (0.64)
R1-zero 32.7 (0.70)
R1-zero-Div (Ours) 34.3 (0.79)
SimpleRL-Zoo 33.9 (0.56)
Eurus-2-7B-PRIME 34.0 (0.74)
E.4 Solution diversity in R1-zero-Div